Circular No. 6140

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

NOVA CENTAURI 1995 R. H. McNaught, Anglo-Australian Observatory, communicates the

following precise position by G. J. Garradd (Loomberah, N.S.W., 0.25-m reflector): α = 13^{h}02^{m}31^{s}.99, δ = −60^{o}11′38′′.1 (equinox 2000.0, uncertainty 0′′.3 in each coordinate). On the measured CCD frame the star is not blended with the GSC star mentioned by W. Liller. P. M. Kilmartin, Mt. John Observatory, communicates a similar position (end figures 31^{s}.85, 38′′.2; residuals of six PPM stars average 0′′.2), measured by her from an astrograph film taken on Feb. 25.49 UT by A. C. Gilmore, who also provides further photometry (cf. IAUC 6139): Feb. 25.62, V = 8.30, U −B = −0.70, B − V = +0.28, V − R = +1.09, V − I = +1.23 (airmass 1.04).

SUPERNOVA 1995G IN NGC 1643 McNaught also reports the following precise position, measured by him

from a 10-min exposure by C. P. Cass with the U.K. Schmidt telescope: α = 4^{h}43^{m}44^{s}.22, δ = −5^{o}18′53′′.8 (equinox 2000.0, PPM, uncertainty 0′′.3 in each coordinate). The offsets from the galaxy center are 4′′.5 east, 16′′.1 north. There is a nearby star of mag ∼ 17 at end figures 43^{s}.50, 10′′.3.

SUPERNOVA 1995E IN NGC 2441 P. Molaro and G. Vladilo, Trieste Observatory; and N. A. Walton,

Royal Greenwich Observatory, La Palma, report that spectrograms (range 350–870 nm, resolution FWHM ∼ 0.6 nm) obtained with the 4.2-m William Herschel Telescope (+ ISIS spectrograph) on Feb. 23.7 UT confirm this object to be a type Ia supernova. The spectrum strongly resembles that of SN 1981B at maximum. Major absorption features are Ca II (857.9 nm), O I (777.3 nm), Mg (448.1 nm) and Si II (635.5, 505.1, 412.9 nm), observed at expansion velocities between −10 000 and −13 000 km/s.

COMET 6P/D’ARREST CCD observations (no flat fielding) by L. H. Wasserman, R. L. Millis

and D. G. Schleicher with the 1.1-m Hall telescope at the Lowell Observatory: Feb. 2.53 UT, R = 21.5, I = 21.9; 3.50, I = 23.

COMET 41P/TUTTLE-GIACOBINI-KRESÁK CCD R magnitudes by S. M. Larson and C. W. Hergenrother with

the Steward Observatory’s 2.3-m reflector, Kitt Peak: Jan. 28.19 UT, 21.2; 28.20, 21.8; 29.28, 21.9.

1995 February 25 Brian G. Marsden